OpenStudy (anonymous):

In the expansion of (2x + a)^3, the co-efficient of the x^2 term is -36. So a must equal A -6 B 6 C -9 D 9 E -3 Show me working out thanks.

OpenStudy (ash2326):

We know that \[(a+b)^3= a^3+3a^2b+3ab^2+b^3\] so her \[(2x+a)^3=8x^3+12x^2a+6xa^2+a^3\] coefficient of x^2 \[(2x+a)^3=8x^3+\underline{12x^2a}+6xa^2+a^3\] 12a it's given 12a=-36 so a=-3

OpenStudy (anonymous):

thanks.